A music industry researcher wants to estimate, with a 90% confidence level, the proportion of young urban people (ages 21 to 35 years) who go to at least 3 concerts a year. Previous studies show that 21% of those people (21 to 35 year olds) interviewed go to at least 3 concerts a year. The researcher wants to be accurate within 1% of the true proportion. Find the minimum sample size necessary.
a) 4142
b) 663
c) 945
d) 4489

Answer by Boreal(15235)

The half interval, 0.01=z(0.95)*sqrt (0.21*0.79)/n
square both sides
0.0001=1.645^2*0.21*0.79/n
n=3.8416*0.21*0.79/0.0001=4489.29, rounds to 4490.
D